OpenStudy (decentnabeel):

A chord is 9.0 cm long and is 3.7cm from the center of the circle.What is the length of the radius to the nearest tenth of a centimeter?

OpenStudy (decentnabeel):

@sweetburger

OpenStudy (decentnabeel):

@rebeccaxhawaii

OpenStudy (decentnabeel):

@AloneS

OpenStudy (kendricklamar2014):



OpenStudy (kendricklamar2014):

\[r^2 = 4.5^2 + 3.7^2\] \[r^2 = 20.25 + 13.69\] \[r^2 = 33.94\] \[r = 5.8 ~c.m.\]

OpenStudy (kendricklamar2014):

Hope it helps @DecentNabeel

OpenStudy (decentnabeel):

yup...i will done

OpenStudy (decentnabeel):

click > http://prntscr.com/axcloz

OpenStudy (kendricklamar2014):

Yep, those are the steps to solve this problem. Nice job :)

OpenStudy (decentnabeel):

Thanks

OpenStudy (kendricklamar2014):

Your welcome